You built a circuit consisting of three parallel plate capacitors with capacitances of 2.0 F, 3.0 F, and 5.0 F, respectively, which are connected in parallel. The whole circuit is connected to a 100 V line. What is the (a) total capacitance, (b) charge stored in each capacitor, and (c) potential difference across each capacitor?



1
Expert's answer
2022-02-15T13:35:03-0500

Given:

"C_1=2.0\:\rm F"

"C_2=3.0\:\rm F"

"C_3=5.0\:\rm F"

"V=100\:\rm V"


(a)  the total capacitance

"C=C_1+C_2+C_3=\rm 2.0+3.0+5.0=10\rm \: F"

(b) charge stored in each capacitor

"{q_1=C_1V=\rm 2.0*100=200\:\rm C}
\\ {q_2=C_2V=\rm 3.0*100=300\:\rm C}
\\{q_3=C_3V=\rm 5.0*100=500\:\rm C}"

(c) potential difference across each capacitor

"V_1=V_2=V_3=V=\rm 100\: V"